# A Novel Pathology Foundation Model by Mayo Clinic, Charité, and Aignostics

### v1

@@ A.2 Additional Results
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Ours Molecular-related HEST-COAD 25.6 ±2.3$\pm$ 2.3± 2.3 25.6 ±4.7$\pm$ 4.7± 4.7 30.7 ±0.0$\pm$ 0.0± 0.0 31.0 ±2.0$\pm$ 2.0± 2.0 25.6 ±3.2$\pm$ 3.2± 3.2 30.6 ±0.3$\pm$ 0.3± 0.3 29.4 ±1.5$\pm$ 1.5± 1.5 HEST-HCC 7.8 ±1.2$\pm$ 1.2± 1.2 8.3 ±0.5$\pm$ 0.5± 0.5 6.8 ±0.4$\pm$ 0.4± 0.4 9.4 ±1.7$\pm$ 1.7± 1.7 9.6 ±1.0$\pm$ 1.0± 1.0 8.4 ±1.2$\pm$ 1.2± 1.2 10.7 ±1.9$\pm$ 1.9± 1.9 HEST-IDC 56.6 ±7.8$\pm$ 7.8± 7.8 58.5 ±7.7$\pm$ 7.7± 7.7 56.8 ±7.6$\pm$ 7.6± 7.6 57.4 ±8.5$\pm$ 8.5± 8.5 59.3 ±8.5$\pm$ 8.5± 8.5 61.0 ±8.1$\pm$ 8.1± 8.1 60.4 ±8.3$\pm$ 8.3± 8.3 HEST-LUAD 54.8 ±2.3$\pm$ 2.3± 2.3 55.2 ±2.2$\pm$ 2.2± 2.2 55.8 ±2.9$\pm$ 2.9± 2.9 57.7 ±1.8$\pm$ 1.8± 1.8 56.9 ±1.7$\pm$ 1.7± 1.7 57.3 ±2.7$\pm$ 2.7± 2.7 58.0 ±1.5$\pm$ 1.5± 1.5 HEST-LYMPH_IDC 24.8 ±4.9$\pm$ 4.9± 4.9 25.8 ±4.1$\pm$ 4.1± 4.1 25.1 ±4.2$\pm$ 4.2± 4.2 25.6 ±3.3$\pm$ 3.3± 3.3 25.9 ±3.3$\pm$ 3.3± 3.3 26.8 ±4.0$\pm$ 4.0± 4.0 26.4 ±4.4$\pm$ 4.4± 4.4 HEST-PAAD 47.9 ±7.0$\pm$ 7.0± 7.0 48.8 ±5.8$\pm$ 5.8± 5.8 49.5 ±5.8$\pm$ 5.8± 5.8 51.1 ±7.8$\pm$ 7.8± 7.8 47.3 ±6.9$\pm$ 6.9± 6.9 50.9 ±4.3$\pm$ 4.3± 4.3 51.8 ±7.4$\pm$ 7.4± 7.4 HEST-PRAD 37.7 ±0.1$\pm$ 0.1± 0.1 32.2 ±8.1$\pm$ 8.1± 8.1 38.4 ±3.1$\pm$ 3.1± 3.1 37.7 ±2.6$\pm$ 2.6± 2.6 35.1 ±3.4$\pm$ 3.4± 3.4 36.1 ±2.0$\pm$ 2.0± 2.0 38.4 ±0.7$\pm$ 0.7± 0.7 HEST-READ 18.5 ±5.9$\pm$ 5.9± 5.9 17.4 ±6.3$\pm$ 6.3± 6.3 18.8 ±6.3$\pm$ 6.3± 6.3 19.9 ±6.8$\pm$ 6.8± 6.8 21.1 ±5.0$\pm$ 5.0± 5.0 24.1 ±2.7$\pm$ 2.7± 2.7 22.8 ±3.1$\pm$ 3.1± 3.1 HEST-SKCM 58.4 ±6.2$\pm$ 6.2± 6.2 63.0 ±2.4$\pm$ 2.4± 2.4 58.8 ±5.8$\pm$ 5.8± 5.8 61.8 ±4.6$\pm$ 4.6± 4.6 63.7 ±3.1$\pm$ 3.1± 3.1 66.1 ±5.8$\pm$ 5.8± 5.8 62.5 ±2.4$\pm$ 2.4± 2.4 HEST-ccRCC 27.3 ±4.0$\pm$ 4.0± 4.0 25.3 ±3.8$\pm$ 3.8± 3.8 24.9 ±4.1$\pm$ 4.1± 4.1 25.3 ±5.2$\pm$ 5.2± 5.2 27.4 ±5.4$\pm$ 5.4± 5.4 29.0 ±3.8$\pm$ 3.8± 3.8 29.4 ±5.5$\pm$ 5.5± 5.5 MSI CRC (patch) 68.8 ±0.1$\pm$ 0.1± 0.1 69.5 ±0.0$\pm$ 0.0± 0.0 70.4 ±0.1$\pm$ 0.1± 0.1 69.9 ±0.1$\pm$ 0.1± 0.1 74.0 ±0.0$\pm$ 0.0± 0.0 71.2 ±0.1$\pm$ 0.1± 0.1 73.6 ±0.0$\pm$ 0.0± 0.0 MSI STAD (patch) 71.2 ±0.1$\pm$ 0.1± 0.1 70.5 ±0.0$\pm$ 0.0± 0.0 71.0 ±0.1$\pm$ 0.1± 0.1 74.1 ±0.1$\pm$ 0.1± 0.1 74.8 ±0.2$\pm$ 0.2± 0.2 73.6 ±0.0$\pm$ 0.0± 0.0 76.0 ±0.1$\pm$ 0.1± 0.1 Molecular-Average 41.6 41.7 42.3 43.4 43.4 44.6 44.9 Morphology-related Pan-cancer TIL 92.9 ±0.0$\pm$ 0.0± 0.0 92.6 ±0.0$\pm$ 0.0± 0.0 92.3 ±0.0$\pm$ 0.0± 0.0 92.6 ±0.1$\pm$ 0.1± 0.1 93.1 ±0.1$\pm$ 0.1± 0.1 93.0 ±0.1$\pm$ 0.1± 0.1 93.0 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (10x) 64.0 ±0.0$\pm$ 0.0± 0.0 68.6 ±0.0$\pm$ 0.0± 0.0 69.1 ±0.0$\pm$ 0.0± 0.0 70.6 ±0.0$\pm$ 0.0± 0.0 73.0 ±0.0$\pm$ 0.0± 0.0 70.4 ±0.0$\pm$ 0.0± 0.0 71.7 ±0.0$\pm$ 0.0± 0.0 TCGA Uniform (20x) 69.4 ±0.0$\pm$ 0.0± 0.0 67.8 ±0.0$\pm$ 0.0± 0.0 68.0 ±0.0$\pm$ 0.0± 0.0 78.1 ±0.0$\pm$ 0.0± 0.0 71.5 ±0.0$\pm$ 0.0± 0.0 72.4 ±0.0$\pm$ 0.0± 0.0 67.8 ±0.0$\pm$ 0.0± 0.0 BACH 73.8 ±0.3$\pm$ 0.3± 0.3 80.1 ±1.0$\pm$ 1.0± 1.0 80.2 ±0.3$\pm$ 0.3± 0.3 76.9 ±0.3$\pm$ 0.3± 0.3 88.7 ±0.3$\pm$ 0.3± 0.3 74.2 ±0.5$\pm$ 0.5± 0.5 92.5 ±0.5$\pm$ 0.5± 0.5 CRC-100k 95.5 ±0.0$\pm$ 0.0± 0.0 95.4 ±0.2$\pm$ 0.2± 0.2 95.9 ±0.0$\pm$ 0.0± 0.0 96.0 ±0.1$\pm$ 0.1± 0.1 96.7 ±0.1$\pm$ 0.1± 0.1 96.2 ±0.1$\pm$ 0.1± 0.1 97.1 ±0.1$\pm$ 0.1± 0.1 MHIST 78.4 ±0.4$\pm$ 0.4± 0.4 84.0 ±0.1$\pm$ 0.1± 0.1 83.1 ±0.1$\pm$ 0.1± 0.1 80.5 ±0.0$\pm$ 0.0± 0.0 85.9 ±0.1$\pm$ 0.1± 0.1 85.0 ±0.2$\pm$ 0.2± 0.2 86.4 ±0.1$\pm$ 0.1± 0.1 PCAM 90.0 ±0.1$\pm$ 0.1± 0.1 93.6 ±0.1$\pm$ 0.1± 0.1 94.3 ±0.1$\pm$ 0.1± 0.1 94.6 ±0.1$\pm$ 0.1± 0.1 93.9 ±0.1$\pm$ 0.1± 0.1 94.3 ±0.1$\pm$ 0.1± 0.1 94.8 ±0.0$\pm$ 0.0± 0.0 CAMELYON16 79.1 ±1.2$\pm$ 1.2± 1.2 85.0 ±0.5$\pm$ 0.5± 0.5 80.6 ±0.7$\pm$ 0.7± 0.7 76.3 ±0.5$\pm$ 0.5± 0.5 86.0 ±0.3$\pm$ 0.3± 0.3 84.0 ±0.8$\pm$ 0.8± 0.8 86.7 ±0.6$\pm$ 0.6± 0.6 PANDA 65.3 ±0.4$\pm$ 0.4± 0.4 69.0 ±0.5$\pm$ 0.5± 0.5 69.6 ±0.7$\pm$ 0.7± 0.7 69.6 ±0.4$\pm$ 0.4± 0.4 66.4 ±1.1$\pm$ 1.1± 1.1 68.0 ±0.6$\pm$ 0.6± 0.6 70.5 ±0.5$\pm$ 0.5± 0.5 Morphology-Average 78.7 81.8 81.5 81.7 83.9 81.9 84.5 Overall Average 57.5 58.9 59.1 59.8 60.8 60.6 61.9

### v2

@@ A.2 Additional Results
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Atlas Molecular-related HEST-COAD 25.6 ±$\pm 2.3$ 25.6 ±$\pm 4.7$ 30.7 ±$\pm 0.0$ 31.0 ±$\pm 2.0$ 25.6 ±$\pm 3.2$ 30.6 ±$\pm 0.3$ 29.4 ±$\pm 1.5$ HEST-HCC 7.8 ±$\pm 1.2$ 8.3 ±$\pm 0.5$ 6.8 ±$\pm 0.4$ 9.4 ±$\pm 1.7$ 9.6 ±$\pm 1.0$ 8.4 ±$\pm 1.2$ 10.7 ±$\pm 1.9$ HEST-IDC 56.6 ±$\pm 7.8$ 58.5 ±$\pm 7.7$ 56.8 ±$\pm 7.6$ 57.4 ±$\pm 8.5$ 59.3 ±$\pm 8.5$ 61.0 ±$\pm 8.1$ 60.4 ±$\pm 8.3$ HEST-LUAD 54.8 ±$\pm 2.3$ 55.2 ±$\pm 2.2$ 55.8 ±$\pm 2.9$ 57.7 ±$\pm 1.8$ 56.9 ±$\pm 1.7$ 57.3 ±$\pm 2.7$ 58.0 ±$\pm 1.5$ HEST-LYMPH_IDC 24.8 ±$\pm 4.9$ 25.8 ±$\pm 4.1$ 25.1 ±$\pm 4.2$ 25.6 ±$\pm 3.3$ 25.9 ±$\pm 3.3$ 26.8 ±$\pm 4.0$ 26.4 ±$\pm 4.4$ HEST-PAAD 47.9 ±$\pm 7.0$ 48.8 ±$\pm 5.8$ 49.5 ±$\pm 5.8$ 51.1 ±$\pm 7.8$ 47.3 ±$\pm 6.9$ 50.9 ±$\pm 4.3$ 51.8 ±$\pm 7.4$ HEST-PRAD 37.7 ±$\pm 0.1$ 32.2 ±$\pm 8.1$ 38.4 ±$\pm 3.1$ 37.7 ±$\pm 2.6$ 35.1 ±$\pm 3.4$ 36.1 ±$\pm 2.0$ 38.4 ±$\pm 0.7$ HEST-READ 18.5 ±$\pm 5.9$ 17.4 ±$\pm 6.3$ 18.8 ±$\pm 6.3$ 19.9 ±$\pm 6.8$ 21.1 ±$\pm 5.0$ 24.1 ±$\pm 2.7$ 22.8 ±$\pm 3.1$ HEST-SKCM 58.4 ±$\pm 6.2$ 63.0 ±$\pm 2.4$ 58.8 ±$\pm 5.8$ 61.8 ±$\pm 4.6$ 63.7 ±$\pm 3.1$ 66.1 ±$\pm 5.8$ 62.5 ±$\pm 2.4$ HEST-ccRCC 27.3 ±$\pm 4.0$ 25.3 ±$\pm 3.8$ 24.9 ±$\pm 4.1$ 25.3 ±$\pm 5.2$ 27.4 ±$\pm 5.4$ 29.0 ±$\pm 3.8$ 29.4 ±$\pm 5.5$ MSI CRC (patch) 68.8 ±$\pm 0.1$ 69.5 ±$\pm 0.0$ 70.4 ±$\pm 0.1$ 69.9 ±$\pm 0.1$ 74.0 ±$\pm 0.0$ 71.2 ±$\pm 0.1$ 73.6 ±$\pm 0.0$ MSI STAD (patch) 71.2 ±$\pm 0.1$ 70.5 ±$\pm 0.0$ 71.0 ±$\pm 0.1$ 74.1 ±$\pm 0.1$ 74.8 ±$\pm 0.2$ 73.6 ±$\pm 0.0$ 76.0 ±$\pm 0.1$ Molecular-Average 41.6 41.7 42.3 43.4 43.4 44.6 44.9 Morphology-related Pan-cancer TIL 92.9 ±$\pm 0.0$ 92.6 ±$\pm 0.0$ 92.3 ±$\pm 0.0$ 92.6 ±$\pm 0.1$ 93.1 ±$\pm 0.1$ 93.0 ±$\pm 0.1$ 93.0 ±$\pm 0.0$ TCGA Uniform (10x) 64.0 ±$\pm 0.0$ 68.6 ±$\pm 0.0$ 69.1 ±$\pm 0.0$ 70.6 ±$\pm 0.0$ 73.0 ±$\pm 0.0$ 70.4 ±$\pm 0.0$ 71.7 ±$\pm 0.0$ TCGA Uniform (20x) 69.4 ±$\pm 0.0$ 67.8 ±$\pm 0.0$ 68.0 ±$\pm 0.0$ 78.1 ±$\pm 0.0$ 71.5 ±$\pm 0.0$ 72.4 ±$\pm 0.0$ 67.8 ±$\pm 0.0$ BACH 73.8 ±$\pm 0.3$ 80.1 ±$\pm 1.0$ 80.2 ±$\pm 0.3$ 76.9 ±$\pm 0.3$ 88.7 ±$\pm 0.3$ 74.2 ±$\pm 0.5$ 92.5 ±$\pm 0.5$ CRC-100k 95.5 ±$\pm 0.0$ 95.4 ±$\pm 0.2$ 95.9 ±$\pm 0.0$ 96.0 ±$\pm 0.1$ 96.7 ±$\pm 0.1$ 96.2 ±$\pm 0.1$ 97.1 ±$\pm 0.1$ MHIST 78.4 ±$\pm 0.4$ 84.0 ±$\pm 0.1$ 83.1 ±$\pm 0.1$ 80.5 ±$\pm 0.0$ 85.9 ±$\pm 0.1$ 85.0 ±$\pm 0.2$ 86.4 ±$\pm 0.1$ PCAM 90.0 ±$\pm 0.1$ 93.6 ±$\pm 0.1$ 94.3 ±$\pm 0.1$ 94.6 ±$\pm 0.1$ 93.9 ±$\pm 0.1$ 94.3 ±$\pm 0.1$ 94.8 ±$\pm 0.0$ CAMELYON16 79.1 ±$\pm 1.2$ 85.0 ±$\pm 0.5$ 80.6 ±$\pm 0.7$ 76.3 ±$\pm 0.5$ 86.0 ±$\pm 0.3$ 84.0 ±$\pm 0.8$ 86.7 ±$\pm 0.6$ PANDA 65.3 ±$\pm 0.4$ 69.0 ±$\pm 0.5$ 69.6 ±$\pm 0.7$ 69.6 ±$\pm 0.4$ 66.4 ±$\pm 1.1$ 68.0 ±$\pm 0.6$ 70.5 ±$\pm 0.5$ Morphology-Average 78.7 81.8 81.5 81.7 83.9 81.9 84.5 Overall Average 57.5 58.9 59.1 59.8 60.8 60.6 61.9